If the total pressure inside a gas tank holding helium gas, oxygen gas, and methane gas was 673 torr, what would be the partial pressure of methane gas if oxygen gas and helium gas each contributed 255 torr?
Answer #1

Total Pressure = 673 Torr

Partial pressure (Po2) of O2 = Partial Pressure (PHe) of He = 255 Torr

Total Pressure = Po2 + PHe + P methane

Total Pressure = 255 + 255 + P methane

P methane = 673 - (255+255) Torr

P methane = 163 Torr